solve algebraically
x^2-4y^2=16
2y-x=2

Answer by stanbon(75887)

solve algebraically
x^2-4y^2=16
2y-x=2
--------------
x = 2y-2
---
Substitute for "x" and solve for "y":
(2y-2)^2 - 4y^2 = 16
4y^2 - 8y + 4 - 4y^2 = 16
8y = 4-16
8y = -12
y = -3/2
-----
Solve for "x":
x = 2y-2
x = 2(-3/2)-2
x = -5
-----------
Cheers,
Stan H.
------------